At a school 100 children have extra lessons of drawing and singing, 90 children's draw 65 sing. How many children's draw and sing ? And how many children only draw ?

Grade:12th pass

1 Answers

Saurabh Koranglekar
3 years ago
Dear student

d+s-(d intersection s ) = 100

90+65 -(d intersection s ) = 100

(d intersection s ) = 55

d=90-55 = 45

Regards